# Fundamental computational limits of weak learnability in high-dimensional multi-index models

### v2

@@ C.4 $g(z_{1},\ldots,z_{p})=z_{1}^{2}+\operatorname{sign}\left(\prod_{j=1}^{p}z_{j}\right)$
$Z_{\rm out}(y,\bm{0},{\bm{I}})$ $=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}$ (73)
@@
C.5 $g(z_{1},\ldots,z_{p})=\sum_{j=1}^{p}\operatorname{sign}(z_{p})$
@@
Appendix D Further numerical observations and details
Here we give more details about the numerical implementation of State Evolution and AMP. Both approaches require to compute ${\bm{g}}_{\rm out}$. All the examples we implemented are detailed in Appendix C.
The integrals are doing with the quadrature package in Scipy. In order to avoid instabilities we regularize the interval of integration by replacing $\infty$ with $\Lambda$. We typically choose $\Lambda\approx 10$. Similarly, we add $\epsilon\approx 10^{-4}$ to the diagonal of $V$. In State Evolution we need to integrate over functions of ${\bm{g}}_{\rm out}$. We do these integrals using a simple Monte Carlo approach. For Figures 1 and 2 we used a total $72000$ samples, for Figure 4 $7200$. Computing such integrals is the numerical bottleneck. In order to make this part faster we parallelised the MCMC: for each iteration of State Evolution we make every worker in our pool estimate the integral, and then average the estimation and then average among workers. In the cases in which ${{\bm{M}}}=0$ is a fixed point, we initialize ${{\bm{M}}}$ with the empirical overlap of AMP at the beginning of the iteration.
In both the AMP and State Evolution implementation we used some damping: the overlap ${{\bm{M}}}$ or the $\hat{{{\bm{{W}}}}}$ at the new iterations are averaged with the current value, with a weight $\delta$ for the new one, where typically $0.90.6<\delta<$ 0.9. We display the evolution of the overlaps in a typical run of AMP for the model $z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$ in Figure 3. We already displayed the values of the overlap and generalisation error at convergence in Figure 2. We can see how AMP has a saddle-to-saddle dynamic, where the algorithm alternates plateaus for $\mathcal{O}(\log{d})$ ) iterations to fast drops in generalisation error, which are associated with new directions being learned.

### v3

@@ F.4 $g(z_{1},\ldots,z_{p})=z_{1}^{2}+\operatorname{sign}\left(\prod_{j=1}^{p}z_{j}\right)$
$Z_{\rm out}(y,\mathbf{0},{\bm{I}})$ $=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}$ (101)
@@
F.5 $g(z_{1},\ldots,z_{p})=\sum_{j=1}^{p}\operatorname{sign}(z_{p})$
Figure 3: Specialisation as an example of grand staircase with $g\!=\!{\rm sign}(z_{1})\!+\!{\rm sign}(z_{2})\!+\!{\rm sign}(z_{3})$. The direction spanned by $z_{1}\!+\!z_{2}\!+\!z_{3}$ is first learned for any $\alpha>0$. The remaining ones are only learned at the specialisation transition occurring at $\alpha\!\approx\!4.3$. We take care of the symmetries as indicated in S.I. G. Crosses denote AMP runs with $d\!=\!100$ averaged over $72$ seeds. The overlap is shown as a function of the sample complexity. Because of the symmetry the elements of $M$ can take one of two values: one on the diagonal and one outside of it. We display $(M_{11}+M_{22}+M_{33})/3$ in blue and $(M_{12}+M_{13}+M_{23})/3$ in red.
@@ F.5 $g(z_{1},\ldots,z_{p})=\sum_{j=1}^{p}\operatorname{sign}(z_{p})$
Appendix G Further numerical observations
Figure 4: Trajectories of $16$ AMP runs with random initialisation at $\alpha=4$ for $g(z_{1},z_{2})={\rm sign}(z_{1}z_{2})$. The shaded areas represent the error on the mean. We can see that as the dimension increases the algorithm gets increasingly slower
Here we give more details about the numerical implementation of State Evolution and AMP. Both approaches require to compute ${\bm{g}}_{\rm out}$. All the examples we implemented are detailed in S.I sec. F.
The integrals are doing with the quadrature package in Scipy. In order to avoid instabilities we regularize the interval of integration by replacing $\infty$ with $\Lambda$. We typically choose $\Lambda\approx 10$. Similarly, we add $\epsilon\approx 10^{-4}$ to the diagonal of $V$. In State Evolution we need to integrate over functions of ${\bm{g}}_{\rm out}$. We do these integrals using a simple Monte Carlo approach. For Figure 1 we used a total $72000$ samples, for Figure 3 $7200$. Computing such integrals is the numerical bottleneck. In order to make this part faster we parallelised the MCMC: for each iteration of State Evolution we make every worker in our pool estimate the integral, and then average the estimation and then average among workers. In the cases in which ${{\bm{M}}}=0$ is a fixed point, we initialize ${{\bm{M}}}$ with the empirical overlap of AMP at the beginning of the iteration.
In both the AMP and State Evolution implementation we used some damping: the overlap ${{\bm{M}}}$ or the $\hat{{{\bm{{W}}}}}$ at the new iterations are averaged with the current value, with a weight $\delta$ for the new one, where typically $0.90.6<\delta<$ 0.9. We display the evolution of the overlaps in a typical run of AMP for the model $z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$ in Figure 2. We already displayed the values of the overlap and generalisation error at convergence in Figure 1. We can see how AMP has a saddle-to-saddle dynamic, where the algorithm alternates plateaus for $\mathcal{O}(\log{d})$ ) iterations to fast drops in generalisation error, which are associated with new directions being learned. We probe experimentally the dependence on size in Figure 4. All the error bars present represent the error on the mean. We used the local computing cluster for our experiments. In particular we needed around $40$ CPUs with $72$ cores each and approximately $3$ Gb per core.